# A Survey on Neural Topic Models: Methods, Applications, and Challenges

### v1

@@ Abstract
Topic models have been prevalent for decades to discover latent topics and infer topic proportions of documents in an unsupervised fashion. They have been widely used in various applications like text analysis and context recommendation. Recently, the rise of neural networks has facilitated the emergence of a new research field—Neural Topic Models (NTMs). Different from conventional topic models, NTMs directly optimize parameters without requiring model-specific derivations. This endows NTMs with better scalability and flexibility, resulting in significant research attention and plentiful new methods and applications. In this paper, we present a comprehensive survey on neural topic models concerning methods, applications, and challenges. Specifically, we systematically organize current NTM methods according to their network structures and introduce the NTMs for various scenarios like short texts and cross-lingual documents. We also discuss a wide range of popular applications built on NTMs. Finally, we highlight the challenges confronted by NTMs to inspire future research.
@@ A Survey on Neural Topic Models: Methods, Applications, and Challenges
Xiaobao ${}^{\ddagger}$,  Thong ${}^{\dagger}$,  Anh Tuan ${}^{\ddagger}$$^{\ddagger}$anyang Technological University,  ${}^{\dagger}$ational University of Singapore xiaobao002@e.ntu.edu.sg,  e0998147@u.nus.edu,  anhtuan.luu@ntu.edu.sg
@@ 1 Introduction
Topic models seek to discover a set of latent topics from a collection of documents, depending on word co-occurrence information. Each topic represents an interpretable semantic concept and is described as a group of related words. For example, a topic about “sports” may relate to words like “baseball”, “basketball”, and “football”. Topic models also infer what topics a document contains (topic proportions) to reveal their underlying semantics. Due to their effectiveness and interpretability, topic models have derived various downstream applications, such as document retrieval, content recommendation, opinion/event mining, and trend analysis (Blei and Lafferty, 2006b; Wang and Blei, 2011; Boyd-Graber et al., 2017; Duong et al., 2022; Churchill and Singh, 2022).
@@
Figure 1: The overview of this survey, including NTMs with different structures, NTMs for various scenarios, applications, and challenges.
@@
Conventional approaches to topic modeling embrace either probabilistic graphical models or non-negative matrix factorization. Approaches based on probabilistic graphical models, such as the classic Latent Dirichlet Allocation (LDA, Blei et al., 2003), have been extensively explored for the past two decades (Blei, 2012). They mainly model the document generation process with topics as latent variables, and then infer model parameters through Variational Inference (Blei et al., 2017) or Monte Carlo Markov Chain (MCMC) methods like Gibbs sampling (Steyvers and Griffiths, 2007). Alternatively, another conventional type of topic models uses non-negative matrix factorization. They directly discover topics by decomposing a term-document matrix into two low-rank factor matrices: one represents words and the other documents (Lee and Seung, 2000; Kim et al., 2015; Shi et al., 2018). These conventional topic models have derived various model structures, such as supervised LDA (Mcauliffe and Blei, 2007) and correlated LDA (Blei and Lafferty, 2006a). Besides the basic topic modeling scenario, researchers have extended topic models to other diverse scenarios, e.g., short text (Yan et al., 2013; Yin and Wang, 2014), cross-lingual (Mimno et al., 2009), and dynamic topic modeling (Blei and Lafferty, 2006b; Wang et al., 2008).
@@
However, despite the achievements of these conventional methods, they generally confront two limitations: (i) Inefficient and labor-intensive parameter inference. These methods necessitate complicated model-specific derivations for parameter inference, and the inference complexity grows along with model complexity. Consequently, this requirement weakens their generalization ability to diverse model structures and application scenarios. (i) Limited scalability to large datasets. Their inference algorithms typically are not parallel, leading to significant time consumption. For example, training a probabilistic dynamic topic model DTM Blei and Lafferty (2006b) on a dataset with 10k documents takes two days (Dieng et al., 2019). Admittedly some parallel inference algorithms have been proposed (Newman et al., 2009; Wang et al., 2009; Liu et al., 2011), but unfortunately they cannot straightforwardly fit other model structures and application scenarios. As a result, how to design effective, flexible, efficient, and scalable topic models has become an urgent imperative in the research field.
@@
In this paper to address these limitations, we present an extensive and up-to-date survey of NTMs, which offers an in-depth and self-contained understanding of NTMs in terms of methods, applications, and challenges. We begin by systematically organizing existing NTMs according to their neural network structures, such as using embeddings or graph neural networks. We then introduce the NTMs designed for various prevalent topic modeling scenarios, e.g., short text, cross-lingual, and dynamic topic modeling, covering a wider range than the early survey (Zhao et al., 2021a). Moreover while omitted by the previous survey (Zhao et al., 2021a), we also organize and discuss the popular applications based on NTMs, developed for diverse tasks like text analysis and text generation. Finally we summarize the key challenges for NTMs in detail to motivate future research directions. Figure 1 depicts the overview of our survey. We conclude the main contributions of this paper as follows:
@@
We accompany this survey with a repository  [^1] https://github.com/bobxwu/Paper-Neural-Topic-Models of the mentioned paper resources to provide easy access for researchers.
@@ 2.1 Problem Setting and Notations
We introduce the problem setting and notations of topic modeling following LDA (Blei et al., 2003). Consider a collection of $N$ documents with $V$ unique words (vocabulary size), and a document is denoted as $\boldsymbol{\mathbf{x}}$. As illustrated in Figure 2, topic models aim to discover $K$ latent topics from this collection. The number of topics $K$ is a hyperparameter, usually determined by researchers manually according to the characteristics of datasets and their target tasks. Each topic is defined as a distribution over the vocabulary, i.e., topic-word distribution, $\boldsymbol{\mathbf{\beta}}_{k}\in\mathbb{R}^{V}$. Then the topic-word distribution matrix of all topics is
@@ 2.2.1 Perplexity
Perplexity, borrowed from language models, measures how a model can predict new documents. It is measured as the normalized log-likelihood of held-out test documents. Perplexity has been used for years to evaluate topic models. Nevertheless, prior studies have empirically demonstrated that perplexity inaccurately reflects the quality of discovered topics as it often contradicts human judgment (Chang et al., 2009). Furthermore, computing log-likelihood is inconsistent among different topic models. This is because they apply various sampling or approximation techniques (Wallach et al., 2009; Buntine, 2009) as well as diverse modeling approaches for topic-word distributions and doc-topic distributions. For instance, certain methods normalize topic-word distributions with respect to topics, some with respect to words, and others opt to keep them unnormalized. These disparities bring challenges to equitable comparisons. Finally, perplexity may not evaluate the practical utility of topic models since users typically employ topic models for content analysis rather than generating new documents (Zhao et al., 2021a; Hoyle et al., 2022). Due to these reasons, perplexity has waned in popularity for topic model evaluation in the recent research field.
@@ 2.2.2 Topic Coherence
Rather than predictive abilities, researchers turn to evaluating the quality of produced topics. For this purpose, researchers propose topic coherence to measure the coherence among the most related words of topics, i.e., top words (determined by topic-word probabilities). Experiments showcase that topic coherence can agree with the human evaluation on topic interpretability (Lau et al., 2014). For example, one widely-used coherence metric is Normalized Point-wise Mutual Information (NPMI, Bouma, 2009; Newman et al., 2010; Lau et al., 2014) [^2] https://github.com/jhlau/topic_interpretability Specifically, the NPMI score between two words $(x_{i},x_{j})$ is calculated as follows:
@@
Later, Röder et al. (2015) propose a new metric, $C_{V}$, which calculates the cosine similarity between NPMI score vectors (Krasnashchok and Jouili, 2018). Given the top $T$ words of a topic, $(x_{1},x_{2},\dots,x_{T})$, the exact calculation of $C_{V}$ is formulated as
@@
The NPMI score computation follows Eq. 2. Röder et al. (2015) empirically demonstrate that $C_{V}$ outperforms previous coherence metrics, NPMI, UCI, and UMass (Mimno et al., 2011), since $C_{V}$ is more consistent with human judgment (See Röder et al. (2015) for experimental results).
@@ 2.2.3 Topic Diversity
These metrics all measure topic diversity based on the uniqueness of individual words. They posit that diversity is optimal when all topics are characterized by distinct top words. However, we question these diversity metrics because of the fact that certain topics naturally share the same words. For example, the word “chip” could be shared by the topics of “potato chip” and “electronic chip”; Similarly, the word “apple” may be covered by the topics of “fruit” and “company”. This issue remains unresolved for reliable diversity evaluation. We in this paper propose a new diversity metric to address this issue (See details in Sec. 7).
@@ 2.2.4 Downstream Task Performance
The performance can be evaluated by accuracy or F1 scores. For document clustering, the common way is to use the most significant topic in a doc-topic distribution as the clustering assignment of a document. Another way is to apply clustering algorithms, e.g., K-Means or DBSCAN, on doc-topic distributions (Zhao et al., 2021b). The clustering performance can be measured by Purity and Normalized Mutual Information (NMI, Manning et al., 2008).
@@ 2.2.5 Visualization
Finally, researchers visualize topic models for evaluation. The typical visualization method is to show the top words of topics and doc-topic distributions, such as using pyLDAvis  [^4] https://github.com/bmabey/pyLDAvis (Sievert and Shirley, 2014) or word cloud  [^5] https://github.com/amueller/word_cloud. Another strategy is to cluster documents on a 2D canvas by reducing the dimension of doc-topic distributions with tools like t-SNE (van der Maaten and Hinton, 2008).
@@ 2.3 Basic NTM based on VAE
We introduce the most basic and popular NTM based on the Variational AutoEncoder (VAE) framework with the neural variational inference technique (Miao et al., 2016; Srivastava and Sutton, 2017). As illustrated in Figure 3, a VAE-based NTM mainly contains an encoder (inference network) and a decoder (generation network). The encoder is to infer doc-topic distributions from documents. To be specific, we use a latent variable $\boldsymbol{\mathbf{r}}\in\mathbb{R}^{K}$ following a logistic normal prior
@@
where $\boldsymbol{\mathbf{\mu}}_{0}$ and $\boldsymbol{\mathbf{\Sigma}}_{0}$ are the mean vector and diagonal covariance matrix respectively. Here the prior distribution is specified with Laplace approximation (Hennig et al., 2012) to approximate a symmetric Dirichlet prior as $\mu_{0,k}=0$ and $\Sigma_{0,kk}=(K-1)/\alpha K$ with hyperparameter $\alpha$ (Srivastava and Sutton, 2017). The variational distribution is modeled by parameters $\Theta$ as
@@
where $\Theta=\{\Theta_{1},\Theta_{2}\}$ and $\mathrm{diag}(\cdot)$ denotes transforming a vector to a diagonal matrix. In practice, we employ MLPs as the encoder networks and transform document $\boldsymbol{\mathbf{x}}$ into a Bag-of-Words (BoW) vector as inputs. Then to avoid gradient variance (Kingma and Welling, 2014; Rezende et al., 2014), we sample $\boldsymbol{\mathbf{r}}$ through the reparameterization trick by sampling a random variable $\boldsymbol{\mathbf{\epsilon}}$:
@@
The first term is the negative expected log-likelihood, i.e., the reconstruction error, where $p_{\Phi}(\boldsymbol{\mathbf{x}}|\boldsymbol{\mathbf{\theta}})$ denotes the generation probability of $\boldsymbol{\mathbf{x}}$. As we sample words from the multinomial distribution, the first term becomes $-\boldsymbol{\mathbf{x}}^{\top}\log(f_{\Phi}(\boldsymbol{\mathbf{\theta}}))$. The second term is the Kullback-Leibler (KL) divergence between the variational and prior distributions, which can be computed through an analytical form (Srivastava and Sutton, 2017). It is also known as a regularization term.
The above is the fundamental structure of a VAE-based NTM. Based on this, NTMs with different structures are proposed to further improve performance and deal with various application scenarios.
@@ 3.2 NTMs with Embeddings
Here $\boldsymbol{\mathbf{W}}\in\mathbb{R}^{D\times V}$ denotes $V$ word embeddings, and $\boldsymbol{\mathbf{T}}\in\mathbb{R}^{D\times K}$ denotes $K$ topic embeddings, where $D$ is the dimension of embedding space. Then Dieng et al. (2020) follow this setting and facilitate topic learning by initializing $\boldsymbol{\mathbf{W}}$ with pre-trained word embeddings like Word2Vec (Mikolov et al., 2013) or GloVe (Pennington et al., 2014). This approach also confers flexibility and efficiency to other topic modeling scenarios. For instance, it is much cheaper to repeat topic embeddings for each time slice in dynamic topic modeling than repeating the entire topic-word distribution matrix (Dieng et al., 2019).
Alternatively, Zhao et al. (2021b) also model topics as embeddings, but use the optimal transport distance between doc-topic distributions and input documents to measure the reconstruction error. Wang et al. (2022a) share the same idea and instead use conditional transport distance. Duan et al. (2022) learn a group of global topic embeddings for task-specific adaptations. Xu et al. (2022) use hyperbolic embeddings to model topics. Due to the tree-likeness property of hyperbolic space, they can capture the hierarchy among topics.
Differently, Wu et al. (2023b) model the topic-word distribution matrix as
@@
Here $\beta_{jk}$ denotes the correlation between $j$-th word and $k$-th topic with $\tau$ as a temperature hyperparameter; $\boldsymbol{\mathbf{w}}_{j}$ is the $j$-th word embedding in $\boldsymbol{\mathbf{W}}$, and $\boldsymbol{\mathbf{t}}_{k}$ is the $k$-th topic embedding in $\boldsymbol{\mathbf{T}}$. Here it computes the Euclidean distance between topic and word embeddings and normalizes overall topics in a softmax manner. This works together with a regularization method in Wu et al. (2023b) to avoid topic collapsing.

### v2

@@ Abstract
Topic models have been prevalent for decades to discover latent topics and infer topic proportions of documents in an unsupervised fashion. They have been widely used in various applications like text analysis and context recommendation. Recently, the rise of neural networks has facilitated the emergence of a new research field—Neural Topic Models (NTMs). Different from conventional topic models, NTMs directly optimize parameters without requiring model-specific derivations. This endows NTMs with better scalability and flexibility, resulting in significant research attention and plentiful new methods and applications. In this paper, we present a comprehensive survey on neural topic models concerning methods, applications, and challenges. Specifically, we systematically organize current NTM methods according to their network structures and introduce the NTMs for various scenarios like short texts and bilingual documents. We also discuss a wide range of popular applications built on NTMs. Finally, we highlight the challenges confronted by NTMs to inspire future research. We accompany this survey with a repository for easier access to the mentioned paper resources: https://github.com/bobxwu/Paper-Neural-Topic-Models.
@@ A Survey on Neural Topic Models: Methods, Applications, and Challenges
Xiaobao Wu   Thong Nguyen   Anh Tuan Luu Nanyang Technological University,  National University of Singapore xiaobao002@e.ntu.edu.sge0998147@u.nus.eduanhtuan.luu@ntu.edu.sg
@@ 1 Introduction
Topic models seek to discover a set of latent topics from a collection of documents, depending on word co-occurrence information. Each topic represents an interpretable semantic concept and is described as a group of related words. For example, a topic about “sports” may relate to words like “baseball”, “basketball”, and “football”. Topic models also infer what topics a document contains (topic proportions) to reveal their underlying semantics. Due to their effectiveness and interpretability, topic models have derived various downstream applications, such as document retrieval, content recommendation, opinion/event mining, and trend analysis Blei and Lafferty (2006b); Wang and Blei (2011); Boyd-Graber et al. (2017); Duong et al. (2022); Churchill and Singh (2022).
@@
Figure 1: The overview of this survey: NTMs with different structures, NTMs for various scenarios, applications of NTMs, and challenges of NTMs.
@@
Conventional approaches to topic modeling embrace either probabilistic graphical models or non-negative matrix factorization. Approaches based on probabilistic graphical models, such as the classic Latent Dirichlet Allocation (LDA, Blei et al., 2003), have been extensively explored for the past two decades Blei (2012). They mainly model the document generation process with topics as latent variables, and then infer model parameters through Variational Inference Blei et al. (2017) or Monte Carlo Markov Chain (MCMC) methods like Gibbs sampling Steyvers and Griffiths (2007). Alternatively, another conventional topic model type uses non-negative matrix factorization. They directly discover topics by decomposing a term-document matrix into two low-rank factor matrices: one represents words and the other documents Lee and Seung (2000); Kim et al. (2015); Shi et al. (2018). These conventional topic models have derived various model structures, such as supervised LDA Mcauliffe and Blei (2007) and correlated LDA Blei and Lafferty (2006a). Besides the basic topic modeling scenario, researchers have extended topic models to other diverse scenarios, e.g., short text Yan et al. (2013); Yin and Wang (2014), cross-lingual Mimno et al. (2009), and dynamic topic modeling Blei and Lafferty (2006b); Wang et al. (2008).
@@
However, despite the achievements of these conventional methods, they generally confront two limitations: (i) Inefficient and labor-intensive parameter inference. These methods necessitate complicated model-specific derivations for parameter inference, and the inference complexity grows along with model complexity. Consequently, this requirement weakens their generalization ability to diverse model structures and application scenarios. (i) Limited scalability to large datasets. Their inference algorithms typically are not parallel, leading to significant time consumption. For example, training a probabilistic dynamic topic model DTM Blei and Lafferty (2006b) on a dataset with 10k documents takes two days Dieng et al. (2019). Admittedly some parallel inference algorithms have been proposed Newman et al. (2009); Wang et al. (2009); Liu et al. (2011), but unfortunately they cannot straightforwardly fit other model structures and application scenarios. As a result, how to design effective, flexible, efficient, and scalable topic models has become an urgent imperative in the research field.
@@
In this paper to address these limitations, we present an extensive and up-to-date survey of NTMs, which offers an in-depth and self-contained understanding of NTMs in terms of methods, applications, and challenges. We begin by systematically organizing existing NTMs according to their neural network structures, such as using embeddings or graph neural networks. We then introduce the NTMs designed for various prevalent topic modeling scenarios, e.g., short text, cross-lingual, and dynamic topic modeling, covering a wider range than the early survey Zhao et al. (2021a). Moreover while omitted by the previous survey Zhao et al. (2021a), we also organize and discuss the popular applications based on NTMs, developed for diverse tasks like text analysis and text generation. Finally we summarize the key challenges for NTMs in detail to motivate future research directions. Figure 1 depicts the overview of our survey. We conclude the main contributions of this paper as follows:
@@ 2.1 Problem Setting and Notations
We introduce the problem setting and notations of topic modeling following LDA Blei et al. (2003). Consider a collection of $N$ documents with $V$ unique words (vocabulary size), and a document is denoted as $\boldsymbol{\mathbf{x}}$. As illustrated in Figure 2, topic models aim to discover $K$ latent topics from this collection. The number of topics $K$ is a hyperparameter, usually determined by researchers manually according to the characteristics of datasets and their target tasks. Each topic is defined as a distribution over the vocabulary, i.e., topic-word distribution, $\boldsymbol{\mathbf{\beta}}_{k}\in\mathbb{R}^{V}$. Then the topic-word distribution matrix of all topics is
@@ 2.2.1 Perplexity
Perplexity, borrowed from language models, measures how a model can predict new documents. It is measured as the normalized log-likelihood of held-out test documents. Perplexity has been used for years to evaluate topic models. Nevertheless, prior studies have empirically demonstrated that perplexity inaccurately reflects the quality of discovered topics as it often contradicts human judgment Chang et al. (2009). Furthermore, computing log-likelihood is inconsistent among different topic models. This is because they apply various sampling or approximation techniques Wallach et al. (2009); Buntine (2009) as well as diverse modeling approaches for topic-word distributions and doc-topic distributions. For instance, certain methods normalize topic-word distributions with respect to topics, some with respect to words, and others opt to keep them unnormalized. These disparities bring challenges to equitable comparisons. Finally, perplexity may not evaluate the practical utility of topic models since users typically employ topic models for content analysis rather than generating new documents Zhao et al. (2021a); Hoyle et al. (2022). Due to these reasons, perplexity has waned in popularity for topic model evaluation in the recent research field.
@@ 2.2.2 Topic Coherence
Rather than predictive abilities, researchers turn to evaluating the quality of produced topics. For this purpose, researchers propose topic coherence to measure the coherence among the most related words of topics, i.e., top words (determined by topic-word probabilities). Experiments showcase that topic coherence can agree with the human evaluation on topic interpretability Lau et al. (2014). For example, one widely-used coherence metric is Normalized Point-wise Mutual Information (NPMI, Bouma, 2009; Newman et al., 2010; Lau et al., 2014) [^1] https://github.com/jhlau/topic_interpretability Specifically, the NPMI score between two words $(x_{i},x_{j})$ is calculated as follows:
@@
Later, Röder et al. (2015) propose a new metric, $C_{V}$, which calculates the cosine similarity between NPMI score vectors Krasnashchok and Jouili (2018). Given the top $T$ words of a topic, $(x_{1},x_{2},\dots,x_{T})$, the exact calculation of $C_{V}$ is formulated as
@@
The NPMI score computation follows Eq. 2. Röder et al. (2015) empirically demonstrate that $C_{V}$ outperforms previous coherence metrics, NPMI, UCI, and UMass Mimno et al. (2011), since $C_{V}$ is more consistent with human judgment (See Röder et al. (2015) for experimental results).
@@ 2.2.3 Topic Diversity
These metrics all measure topic diversity based on the uniqueness of individual words. They posit that diversity is optimal when all topics are characterized by distinct top words. However, we question these diversity metrics because certain topics naturally share the same words. For example, the word “chip” could be shared by the topics of “potato chip” and “electronic chip”; Similarly, the word “apple” may be covered by the topics of “fruit” and “company”. This issue remains unresolved for reliable diversity evaluation. We in this paper propose a new diversity metric to address this issue (See details in Sec. 7).
@@ 2.2.4 Downstream Task Performance
The performance can be evaluated by accuracy or F1 scores. For document clustering, the common way is to use the most significant topic in a doc-topic distribution as the clustering assignment of a document. Another way is to apply clustering algorithms, e.g., K-Means or DBSCAN, on doc-topic distributions Zhao et al. (2021b). The clustering performance can be measured by Purity and Normalized Mutual Information (NMI, Manning et al., 2008).
@@ 2.2.5 Visualization
Finally, researchers visualize topic models for evaluation. The typical visualization method is to show the top words of topics and doc-topic distributions, such as using pyLDAvis  [^3] https://github.com/bmabey/pyLDAvis Sievert and Shirley (2014) or word cloud  [^4] https://github.com/amueller/word_cloud. Another strategy is to cluster documents on a 2D canvas by reducing the dimension of doc-topic distributions with tools like t-SNE van der Maaten and Hinton (2008).
@@ 2.3 Basic NTM based on VAE
We introduce the most basic and popular NTM based on the Variational AutoEncoder (VAE) framework with the neural variational inference technique Miao et al. (2016); Srivastava and Sutton (2017). As illustrated in Figure 3, a VAE-based NTM mainly contains an encoder (inference network) and a decoder (generation network). The encoder is to infer doc-topic distributions from documents. To be specific, we use a latent variable $\boldsymbol{\mathbf{r}}\in\mathbb{R}^{K}$ following a logistic normal prior
@@
where $\boldsymbol{\mathbf{\mu}}_{0}$ and $\boldsymbol{\mathbf{\Sigma}}_{0}$ are the mean vector and diagonal covariance matrix respectively. Here the prior distribution is specified with Laplace approximation Hennig et al. (2012) to approximate a symmetric Dirichlet prior as $\mu_{0,k}=0$ and $\Sigma_{0,kk}=(K-1)/\alpha K$ with hyperparameter $\alpha$ Srivastava and Sutton (2017). The variational distribution is modeled by parameters $\Theta$ as
@@
where $\Theta=\{\Theta_{1},\Theta_{2}\}$ and $\mathrm{diag}(\cdot)$ denotes transforming a vector to a diagonal matrix. In practice, we employ MLPs as the encoder networks and transform document $\boldsymbol{\mathbf{x}}$ into a Bag-of-Words (BoW) vector as inputs. Then to avoid gradient variance Kingma and Welling (2014); Rezende et al. (2014), we sample $\boldsymbol{\mathbf{r}}$ through the reparameterization trick by sampling a random variable $\boldsymbol{\mathbf{\epsilon}}$:
@@
The first term is the negative expected log-likelihood, i.e., the reconstruction error, where $p_{\Phi}(\boldsymbol{\mathbf{x}}|\boldsymbol{\mathbf{\theta}})$ denotes the generation probability of $\boldsymbol{\mathbf{x}}$. As we sample words from the multinomial distribution, the first term becomes $-\boldsymbol{\mathbf{x}}^{\top}\log(f_{\Phi}(\boldsymbol{\mathbf{\theta}}))$. The second term is the Kullback-Leibler (KL) divergence between the variational and prior distributions, which can be computed through an analytical form Srivastava and Sutton (2017). It is also known as a regularization term.
The above is the fundamental structure of a VAE-based NTM, followed by pioneer studies like NVDM Miao et al. (2016) and ProdLDA Srivastava and Sutton (2017). Based on this, NTMs with different structures are proposed to further improve performance and deal with various application scenarios.
@@ 3.2 NTMs with Embeddings
Here $\boldsymbol{\mathbf{W}}\in\mathbb{R}^{D\times V}$ denotes $V$ word embeddings, and $\boldsymbol{\mathbf{T}}\in\mathbb{R}^{D\times K}$ denotes $K$ topic embeddings, where $D$ is the dimension of embedding space. Then Dieng et al. (2020) follow this setting and propose ETM (Embedding Topic Model). ETM facilitates topic learning by initializing $\boldsymbol{\mathbf{W}}$ with pre-trained word embeddings like Word2Vec Mikolov et al. (2013) or GloVe Pennington et al. (2014). This approach also confers flexibility and efficiency to other topic modeling scenarios. For instance, it is much cheaper to repeat topic embeddings for each time slice in dynamic topic modeling than repeating the entire topic-word distribution matrix Dieng et al. (2019).
Alternatively, Zhao et al. (2021b) propose NSTM. It also models topics as embeddings, but uses the optimal transport distance between doc-topic distributions and input documents to measure the reconstruction error. Wang et al. (2022a) share the same idea and instead use conditional transport distance. Duan et al. (2022) learn a group of global topic embeddings for task-specific adaptations. Xu et al. (2022) propose HyperMiner, using hyperbolic embeddings to model topics. Due to the tree-likeness property of hyperbolic space, they can capture the hierarchy among topics.
Differently, Wu et al. (2023b) propose ECRTM, which models the topic-word distribution matrix as
@@
Here $\beta_{jk}$ denotes the correlation between $j$-th word and $k$-th topic with $\tau$ as a temperature hyperparameter; $\boldsymbol{\mathbf{w}}_{j}$ is the $j$-th word embedding in $\boldsymbol{\mathbf{W}}$, and $\boldsymbol{\mathbf{t}}_{k}$ is the $k$-th topic embedding in $\boldsymbol{\mathbf{T}}$. It computes the Euclidean distance between topic and word embeddings and normalizes overall topics in a softmax manner. This works together with a clustering regularization method. The regularization considers topic embeddings as cluster centers and word embeddings as cluster samples; then it forces topic embeddings to be the centers of separately aggregated word embeddings by optimal transport. This effectively avoids the topic collapsing issue where topics are repetitive to each other.